# A new criterion for interpreting acoustic emission damage signals in condition monitoring based on the distribution of cluster onsets

### v1

@@ 3.2.2 Results
The tests presented below were made using a brand new electrode and after cleaning the machine. Two tests were performed successively. Algorithm 2 (based on onsets) was used on every combination of $3$ features within the set of $24$ features (total of 2024 combinations). The $20$ best partitions found by this algorithm for each number of clusters $K\in\{4,6,8,10\}$ were used to build the histograms.
(a) Test #1.
(b) Test #2.
Figure 10: Histogram of onsets (all peaks below 10 are not represented to lighten the figure).
Histograms are presented in Figure 10 for test #1 (brand new electrode) and test #2 (after one test). In these experiments, the electrode position is not known precisely due to its continuous wear during machining. Therefore, we used the camera to evaluate, approximately, when the first sparkle appears at the opening of the hole into the cavity. Based on the images, we identified the following time instants (depicted on figure 10):
  • Test #1: The first sparkle after opening the hole appears between 14-15 s and the next surface is drilled from about 17-18 s.
  • Test #2: The first sparkle after opening the hole appears at about 18-19 s and the next surface is drilled from about 22-23 s.
The cumulated energy is superimposed on histograms. The energy feature is traditionnally extracted from AE signals since it provides some trends about damage evolution.
For test #1 (Fig. 9(a)), four main stages identified on the histograms are represented by dash-dotted lines. The position of the lines is based on the evolution of the cumulated energy and on the position of onsets of clusters. In stage I (0-6 s), there are three primary peaks and a low energy. They are related to the first appearances of sparkles at the gap distance (defined as the distance between the electrode and the workpiece for which the first sparkles appear). During the process, drilling is performed if a constant gap is ensured between the electrode and part surfaces. However, these surfaces are limited in the first stage, and a few number of sparkles are required until an initial hole is actually made. In Stage II (6-15 s), the cumulated energy increases importantly, in two main phases (6-12 s, 12-15 s) with one primary onset when the slope changes (around 12 s). The changes of the slope in energy during drilling is typical from an adaptation of the machine to find a control law in order to succeed in drilling. At 15 s, Stage III starts, until about 17.6 s, with two primary onsets (number 5 and 6). The rectangle points out this important stage because the beginning of Stage III corresponds to the first sparkle observed on the camera. Therefore, this onset clearly points out that the drilling is approaching the output of the first part of the workpiece. This rectangle shows that about 2.6 s are needed to finalize the drilling and approaching the surface of the second part of the workpiece. From 18 s onwards (Stage IV), we can see one more onset (number 7) representing the start of drilling of the second part. No more significant onsets appear meaning that the AE signals collected in the second part have similar characteristics as in the first part, likewise the machine was using a similar control law.
For test #2 (Fig. 9(b)), the length of Stage IV is shorter because the process was able to drill more easily the second part of the workpiece from about 21.5 s. The process is more stable than in the case of test #1 in order to find its control law to succeed in drilling. This is confirmed by a more regular cumulated energy increase until the first sparkle observed with the camera at the opening of the hole. Indeed, test #1 depicted a significant variation of the cumulated energy curve associated to the start of a stable drilling process. In test #2, the drilling process is quite continuous and stable. It can be observed that the left-hand side of the rectangle points out the important phase when the first sparkle occurs (around 18.2 s). Compared to test #1, there is no onset exactly at the same location, but there is one just before (number 7) and one just after (number 8). The latter (within the rectangle) represents the first sparkle after opening the hole in the cavity, occurring around 19.4 s. The end of this important phase (from an industrial point of view) is characterised by one onset (number 9) just before the start of the drilling of the next part (Stage IV).
3.3 Data set #3: Monitoring of an hybrid composite/metal joint structure
@@
The last application is about monitoring of damages in a hybrid composite/metal joint structure by using the AE technique. The proposed criterion is used to evaluate when damages occur and how it grows during loading until the failure of a metal/composite joint.
@@ 3.3.1 Materials and method
The studied specimen is a composite tube manufactured by UTC Aerospace Systems with a joint between a composite part and a metal part. The composite part comprised eight plies of filament wound carbon fibre tows, used to form a tube with a nominal internal diameter of 45.2 mm and wall thickness of 4 mm. The tube geometry was selected to target loads within the range of the target aerospace applications of the structure where bonded joints are harder to certify.
The joining technology between the metallic end-fitting and the composite tube is in the form of a bespoke helical thread (screw thread). During the assembly of each joint, an epoxy-based adhesive is applied to the metallic screw thread for lubrication purposes. Preloading of the joint results in the composite being subjected to a through-thickness contact pressure and at the interface, loads are transmitted in two ways: (i) through a reaction force at multiple teeth, and (ii) via friction which results from the predominant contact pressure at flat regions within the bespoke helical thread. If the frictional forces are surpassed, relative movement can occur between the composite and metallic parts of the joint. For this reason, it is of critical importance to the designer to consider both slip and matrix damage. The test of the hybrid joint was designed to prioritise the failure of the interface regions, rather than failing within the composite tube or metallic end fitting.
The specimen was tested in a quasi-static tensile loading configuration on an Instron 5989 electro-mechanical testing machine fitted with a 600 kN load cell. It comprises a composite tube and two metallic end fittings that form two joints, connected to the loading rig via rod ends with spherical bearings, which are screwed into the metallic end fittings. These rod ends were then pinned to clevises that are screwed into the rig. The described test set-up is shown in Figure 11-a).
@@
The specimen was loaded at 0.5 mm/min crosshead speed in a single cycle until failure. The specimen failed at the instrumented joint. The breaking loads was 268 kN. The final failure consisted of significant damage caused by initial matrix cracking and followed by slippage at the composite/metal interface of one of the two joints (Figure 11-d) where the screw threads have been almost erased during failure.
@@
AE data were recorded by a PCI-2 based system, and captured in the software "AEWin" (Mistras, USA), with a sampling rate of 10 MHz and 20 dB of pre-amplification per sensor. Hit detection was performed by the software using a threshold of 55 dB and HDT=800 µs, PDT=200 µs and HDT=1 ms. In the following, only the data from sensors on the composite part (number 3, 4 on Fig. 11b)) were used for the analysis. The same 19 features as in the first application were used.
@@ 3.3.2 Results
The goal of clustering with onset-based CVI in this application is to determine whether it is able to highlight some phenomenon in a complex engineering structure and if it can help to detect incipient damages within the joint. In many engineering structures, the use of data-driven methods such as clustering is appealing because it can reveal patterns and associations in the data that can help to identify the damage mechanisms. This is particularly of use where the data set is large but also where our predictions and a priori knowledge have a high uncertainty (like with the metal/composite joint). In these cases, the end-user of the AE technique does not have the confidence to assign labels and cannot run supervised algorithms.
@@
In the following, the Gustafson-Kessel clustering method [32] was used in Algorithm 2 by considering all combinations of subsets of 3 features among 19 (so a total of $969$ combinations), and varying the number of clusters $K\in\{4,6,8,10\}$ leading to $3876$ partitions. The $20$ best partitions found by this algorithm for each number of clusters were used to build the histograms represented in Figure 11(a). Since we do not have access to a ground truth for this data set, we superimposed the cumulated energy as presented in the previous application and we provided the distribution of amplitudes in clusters for $K=10$ (Fig. 11(b)).
(a) Histogram of onsets according to time of test and load.
(b) Distribution of amplitudes in clusters.
Figure 12: Analysis of clustering results for the third application.
Based on these figures, we can extract three main stages before the ruin. Each stage is characterised by a different evolution of the cumulated energy and different distributions of amplitudes. In the first two stages ($0-275$ s and $265-465$ s), the cumulated energy shows two plateaus. Actually, there are AE signals during these two periods but their energy is small compared to the signals observed later on, for example during stage III. This is confirmed by Fig. 11(b) where the first three clusters occurring during stage I depict the lowest amplitudes ($55-75$ dB). Stage II, where onsets $4-7$ are detected, corresponds to middle range amplitudes ($75-100$ dB). Stage III corresponding to the gradual increase of the cumulated energy is characterised by the largest amplitudes ($>100$ db).
The histograms of onsets (Fig. 11(a)) depicts a primary peak at $270$ s. This peak is remarkably close to the first step of increase of the cumulated energy, which corresponds to a sudden release of energetic AE signals. The accumulation of loads during stage I reaches a point that leads to the primary peaks #4 and #5 observed at $265$ s and $305$ s.
During the plateau $275-465$ s, few onsets are observed. Since the load is still increasing, it means that previously detected clusters are accumulating until some highly energetic AE signals occurring around $465$ s and corresponding to a new primary peak (#7). The position of this peak is remarkably close to the beginning of the increase of the cumulated energy corresponding to the starting point of a cascade of damages until failure.
@@
A simulation model was developed using Abaqus [33, Chap. 7]. The estimated slip accumulation was found between 90-180 kN, followed by a fatigue period until ultimate failure at 280 kN (compared to 262 kN experimentally). Therefore, according to Figure 11(a), slippage can correspond to Stage II found by AE, with clusters #4-8. It corroborates the fact that some damages are described by multiple clusters. Slippage in the joint is indeed a complex phenomenon which can not be described by a unique cluster. The high number of onsets in Stage I, before 265 s (90 KN), together with the low cumulated energy observed in this stage, can be related to friction in the joint when the loads are transmitted at the teeth leading to slippage in Stage II. The accumulation of the loads, which appeared in the two previous stages, starts being critical at 465 s (about 160 kN) corresponding to the fatigue period found by simulation (Stage III in AE data). The clusters found around 600 s (corresponding to 78% of the test duration and 205 kN) are incipient critical damages leading to final failure detected with cluster #10.
4 Conclusion and future works
This study proposes a new approach to interpret acoustic emission data from clustering results. Conversely to standard approaches which favor results with compact and separable clusters, our approach is based on a new Clustering Validation Index (CVI) that put attention on how onsets of clusters are distributed on the time/load scales. Onsets reflect the initiation of a “defect”, as illustrated in this study by a loosening of a mechanical assembly, a new stage in a manufacturing process or a damage of a composite structure. In most of monitoring applications, this initiation is a very important characteristic (in the sense that two distinct defects will generally have different onsets). Therefore, focusing on onsets to interpret AE data can help in monitoring the progression of different damages.
On a benchmark data set, the proposed onset-based CVI outperformed the standard approaches in detecting seven levels of loosening of bolted plates under vibration. The criterion also allowed to detect the changes in the loosening with higher accuracy. These results demonstrate the suitability of the criterion to select the best clustering results without a priori knowledge. Moreover, in case of a priori knowledge is available, the proposed criterion is able to incorporate it which makes it unique compared to standard approaches.
In the second application, data were collected during micro-drilling of hard material using Electrical Discharge Machining. The analysis of the data based on the proposed onset-based criterion allowed us to underline different steps in the process. That paves the way for the control of the quality of the machining through a feedback from acoustic emissions.
The third application concerned the analysis of data collected on a hybrid composite/metal joint. The onset-based criterion allowed us to draw a scenario about the damage occurring within the joint until slippage and failure. Discriminating the onset of the slipping phases leading to failure is essential for Structural Health Monitoring of such hybrid materials.
This work illustrates the need to develop specific tools dedicated to AE data analysis in order to better interpret the behavior of the materials or systems during monitoring. The onset-based criterion is one of these tools and we have shown its relevance on three different applications. As future work, we aim at using this criterion in supervised learning algorithms for Structural Health Monitoring purposes. As shown in experiments, the criterion helps in detecting changes in data which can be exploited to accurately detect incipient defects and to trigger prognostics algorithms. We also aim at developing damage models able to generate onsets of acoustic emission through time or load and to use those a priori onsets in the proposed criterion for improving clustering and classification based on physically-informed algorithms.
Acknowledgement
This work was partly carried out in the framework of the EIPHI Graduate school (contract ANR-17-EURE-0002).
@@ 29 $\theta^{*}\leftarrow\theta(S^{*},K^{*})$;
Appendix B Onset-based algorithm
1 Input:        Matrix of features $\mathbf{Y}$ with size $N\times d$;
3 Input:        The subsets of features $\mathcal{S}$;
4 Output:      Histogram of onsets;
5 for $K\in[K_{\textrm{min}},K_{\textrm{max}}]$ do
6       for $S\in\mathcal{S}$ do
8             $[P(S,K),\theta(S,K)]\leftarrow$ Clustering($\mathbf{Y}(:,S)$, $K$);
9             // Compute quality according to how onsets are distributed;
11
12       end for
16       // Find onsets of the selected partitions;
17       $\textrm{Onsets}(K)=\textrm{Find\_onsets}(F(K))$;
18
19 end for

### v3

@@ 3.2.2 Results
The tests described below were conducted using a brand new electrode and after cleaning the machine. Two tests were performed successively. Algorithm 1, which is based on onsets, was applied to every combination of 3 features from a set of 24 (totalling 2024 combinations). The 20 best partitions for each number of clusters $K\in\{4,6,8,10\}$ were used to generate the histograms.
Figure 10: Histogram of onsets (peaks below 10 are not represented to lighten the figure).
Figure 10 presents histograms for Test #1 (with a new electrode) and Test #2 (after one use), showing peaks with more than 10 occurrences. Due to the continuous wear of the electrode during machining, its position was estimated using a camera to determine when the first spark appeared at the hole’s opening. The time instances identified are: 1) Test #1: First spark at 14-15 s; drilling from 17-18 s; Test #2: First spark at 18-19 s; drilling from 22-23 s. Cumulated energy is superimposed on the histograms, as this feature traditionally indicates damage evolution trends.
For Test #1 (Fig. LABEL:edmtest1), four main stages are identified by the dashed lines, based on cumulated energy and onset positions. Stage I (0-6 s) is characterised by three primary peaks with low energy, related to initial spark appearances. Drilling progresses with a few sparkles until a hole is made. Stage II (6-15 s) shows significant increase in cumulated energy in two phases (6-12 s, 12-15 s), with a primary onset around 12 s indicating the machine’s adaptation to drilling. Stage III (15-17.6 s) depicts two primary onsets (5 and 6) and corresponds to the first spark observed by the camera, marking the approach to the first part of the work-piece. Stage IV (18 s onwards) shows one more onset (7) indicating the start of drilling the second part, with similar characteristics to the first part. For Test #2 (Fig. LABEL:edmtest2), Stage IV is shorter, and the process is more stable, starting at about 21.5 s. The cumulated energy increases more regularly, indicating a more stable drilling process. The rectangle highlights the first spark (around 18.2 s), with onsets before (7) and after (8) the observed spark. The end of this phase (just before starting drilling the next part) is marked by onset (9).
3.3 Dataset #3: Hybrid composite/metal joint structure
@@
This application is about monitoring of damages in a hybrid composite/metal joint structure. The proposed criterion is used to evaluate when damage mechanisms occur and how they grow during loading until the failure of the composite/metal joint.
@@ 3.3.1 Materials and method
The specimen studied is a composite tube manufactured by Collins Aerospace, featuring a joint between a composite part and a metal part. The composite section consists of eight plies of filament-wound carbon fibre tows, forming a tube with a nominal internal diameter of 45.2 mm and a wall thickness of 4 mm. This geometry was chosen to handle loads typical of aerospace applications, where certifying bonded joints can be challenging.
The joint technology integrates a bespoke helical screw thread, where an epoxy-based adhesive is applied to the metal thread for lubrication during assembly. Preloading the joint applies through-thickness contact pressure to the composite, and loads are transmitted in two ways: (i) through reaction forces at multiple teeth and (ii) through friction at flat regions within the helical thread. If the frictional forces are exceeded, relative movement between the composite and metal parts can occur. Therefore, it is crucial to consider both slip and matrix damage in the design. The test was designed to ensure failure occurs at the interface regions rather than within the composite tube or metal end fitting.
The specimen was tested in a quasi-static tensile loading configuration on an Instron 5989 electro-mechanical testing machine fitted with a 600 kN load cell. It comprises a composite tube and two metallic end fittings that form two joints, connected to the loading rig via rod ends with spherical bearings, which are screwed into the metallic end fittings. These rod ends are pinned to clevises that are screwed into the rig. The described test set-up is shown in Figure 11-a).
@@
The specimen was loaded at 0.5 mm/min crosshead speed in a single cycle until failure. The specimen failed at the instrumented joint. The breaking load was 268 kN. The final failure consisted of significant damage caused by initial matrix cracking and followed by slippage at the composite/metal interface of one of the two joints (Figure 11-d) where the screw threads have been almost erased during failure.
@@
AE data were recorded by a PCI-2 based system, and captured in the software “AEWin” (Mistras, USA), with a sampling rate of 10 MHz and 20 dB of pre-amplification per sensor. Hit detection was performed by the software using a threshold of 55 dB and HDT=800 µs, PDT=200 µs and HDT=1 ms. In the following, only the data from sensors on the composite part (number 3, 4 on Fig. 11b)) were used for the analysis. The same 19 features as in the first application were used.
@@ 3.3.2 Results
The objective of applying the onset-based CVI in this context is to assess its capability to reveal phenomena in a complex engineering structure and to detect incipient damages within the joint. Data-driven methods like clustering can uncover patterns in large datasets, which is especially useful when prior knowledge is uncertain, as with the metal/composite joint scenario. This approach is beneficial when assigning labels is challenging and supervised algorithms are not feasible.
@@
The Gustafson-Kessel clustering method [21] was employed in Algorithm 1, considering all combinations of 3 features out of 19 (totalling 969 combinations) and varying the number of clusters $K\in\{4,6,8,10\}$, resulting in 3876 partitions. The 20 best partitions for each K were used to construct the histograms shown in Figure LABEL:histoa. Due to the lack of ground truth for this dataset, we compared the results with the cumulated energy from the previous application and provided the distribution of amplitudes in clusters for $K=10$ (Fig. LABEL:histob).
Figure 12: Analysis of clustering results for the third application. Left: Histogram of onsets according to time of test and load. Right: Distribution of amplitudes in clusters.
The analysis reveals three distinct stages before failure. Each stage shows a unique evolution of cumulated energy and amplitude distributions. In the first two stages (0-275 s and 265-465 s), the cumulated energy exhibits two plateaus, indicating lower energy AE signals compared to the later stage III. This is confirmed by Fig. LABEL:histob, where clusters in stage I have the lowest amplitudes (55-75 dB), stage II shows middle-range amplitudes (75-100 dB), and stage III has the highest amplitudes ($>$100 dB).
The histogram of onsets (Fig. LABEL:histoa) shows a primary peak at 270 s, closely aligning with the initial increase in cumulated energy, marking a significant release of AE signals. This peak corresponds to primary peaks #4 and #5 at 265 s and 305 s, indicating incipient damage. During the plateau phase (275-465 s), fewer onsets are observed, with a significant peak (#7) at 465 s, aligning with the start of increased cumulated energy and subsequent damage cascade.
@@
Simulation results using Abaqus [9, Chap. 7] show slip accumulation between 90-180 kN, followed by fatigue until failure at 280 kN (compared to 262 kN experimentally). This correlates with Stage II in the AE data, involving clusters #4-8, suggesting that damages are represented by multiple clusters. The high number of onsets before 265 s (90 kN) and the low cumulated energy during this period may indicate friction, while the critical load accumulation at 465 s (160 kN) aligns with the fatigue period found in simulations (Stage III). Clusters around 600 s (78% of test duration, 205 kN) represent incipient critical damages leading to final failure detected with cluster #10.
4 Conclusion and future work
This study introduces a novel approach for interpreting acoustic emission (AE) data through a new clustering validation index (CVI) that emphasises the distribution of cluster onsets over time, cycles, or load scales. Unlike traditional methods that prioritise compact and separable clusters, our approach focuses on onsets, which are crucial for identifying different defects, such as mechanical loosening, process stages, or composite damage. This shift in focus is particularly valuable in monitoring applications where the initiation of defects is a key characteristic.
The results demonstrated the criterion’s ability to select optimal clustering results without prior knowledge, and to incorporate prior knowledge when available, which underscores its uniqueness and applicability.
In the bolted-joint benchmark, histograms of cluster onsets revealed that the onset-based CVI detected six out of seven true loosening events (missing only the final one), whereas the shape-based scheme missed three and introduced false positives. Quantitatively, the onset-based CVI improved the worst-case Rand index by up to 7% and eliminated all low-quality outliers. This demonstrates that anchoring validation on the temporal distribution of cluster onsets yields more robust and interpretable partitions for SHM applications.
Applying the onset criterion to micro-EDM drilling revealed four distinct operational stages corresponding to tool engagement, stable drilling, hole breakthrough, and secondary penetration. Four main stages were identified by the onset-based CVI. The data-driven insights point to the potential for real-time quality control and tool-wear monitoring in industrial EDM processes.
In the hybrid composite–metal joint tests, onset histograms and amplitude-cluster distributions exposed three damage phases under quasi-static tensile loading: an early friction-dominated slip, a fatigue-driven plateau and a rapid damage cascade leading to catastrophic failure. The primary onset matched simulation-predicted fatigue onset at  280 kN, while a secondary peak marked the transition to rapid interface degradation. These results highlight the criterion’s ability to detect incipient interfacial slip and matrix cracking without prior labeling, offering a powerful tool for complex joint monitoring.
By focusing on when clusters appear rather than solely on their geometric compactness, the onset-based CVI aligns more directly with physical damage progression, making it especially suited for progressive degradation monitoring. However, its performance depends on sufficient temporal resolution and a known or consistent test duration. Adaptive schemes for estimating the duration or windowed onset tracking could be developed. Moreover, small clusters (e.g., very early or late events) can introduce variability, as seen in low-occurrence peaks; integrating amplitude or energy weighting could mitigate this.
Building on the aforementioned findings, future work will (i) integrate the onset-based index into supervised prognostics frameworks to trigger real-time alerts for defect initiation, (ii) explore hybrid CVIs combining onset information with traditional shape metrics for richer interpretability, and (iii) develop physics-informed priors for $p_{\rm{true}}$ to guide clustering in domains where simulations can predict expected damage timelines. Finally, embedding onset tracking within streaming algorithms will extend applicability to continuous SHM deployments.